the sum of 4 term of an A.p is 4 and the sum of the product of the first and the last and the 2 middle term is -3 find the numbers??
Answer by mananth(16949)

let the 4 numbers be
a-3d,a-d,a+d,a+3d
a-3d+a-d+a+d+a+3d=4
4a=4
a=1
(a-d)(a+d)+(a-3d)(a+3d)=-3

a^2-d^2+a^2-9d^2=-3
2a^2-10d^2=-3
but a=1
2-10d^2=-3
10d^2=5
d^2=1/2
d= +/- 1/sqrt(2)
1-(3/sqrt(2)), 1-1/sqrt(2),1+1/sqrt(2), 1+(3/sqrt(2))
plug d= -1/sqrt(2) and get the other 4 numbers